In 2019, Carson is claimed as a dependent on his parents' tax return. Carson's parents provided most of his support. What is Carson's tax liability for the year in each of the following alternative circumstances? Use Tax Rate Schedule, Dividends and Capital Gains Tax Rates, Estates and Trusts for reference. a. Carson is 17 years old at year-end and earned $14,150 from his summer job and part-time job after school. This was his only source of income. (Round your final answer to the nearest whole dollar amount.)
Answer #1

Answer:

According to Internal revenue code individual is reliant and younger than 65 and his earned just income doesn't exclude $12000 then he not required to file return.

Here Carson is 17 year old and dependent on his parents and earned from his summer job and part - time work after school $14,150 which is above $12000 ,so he needs to file return of income.

Standard deduction more prominent of following :

(1) $1050

(2) Earned pay + $350 (yet the all out can't be more than essential standard deduction)

So here $12000

So $12000 is higher accessible as deduction.

  • So $14,150 - $12000 = $ 2,150
  • Tax = $ 2,150 * 10% = $ 2,150 *10/100 = $ 2,150*0.1 = $ 215

Therefore ,Carson's tax liability for the year 2019 is $ 215.